# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 4.4 Marrone-Treanor Pseudo-Temperature
For each condition in the equilibrium QCT dataset, we therefore determine the value of $U$ that matches $\langle\varepsilon_{v,s}\rangle_{d}(T,T_{v}=T)$, via a Newton iteration on Eq. (16). We now have the dependence of $U$ on $T$. This dependence is observed to be fit well by a linear relationship in inverse temperature space, as in Eq. (15). The results are shown in Fig. 6. Because we are interested in minimizing the error in vibrational energy decrease per dissociation, the model parameters $a_{U}$ and $U^{*}$ are determined by fitting to $\langle\varepsilon_{v,s}\rangle_{d}$ directly, by substituting Eq. (15) into Eq. (16). In this fashion, we have computed the dependence of $U$ on $T$ based on the equilibrium vibrational energy change due to dissociation.
@@
Figure 6: Dependence of Marrone-Treanor model parameter $U$ on $T$, determined using the change in vibrational energy per dissociation in the equilibrium test set. The linear fit in inverse temperature space, Eq. (15), is also shown.
@@ 4.5 Comparison of MMT Model with QCT Data
QCT data are compared to the model in Fig. 7 for nitrogen dissociation and Fig. 8 for oxygen dissociation. This modified Marrone-Treanor model is found to accurately fit all quantities of interest for all dissociation reactions.
@@
It is worth noting that all model parameters have been determined using data in the equilibrium test set only: dissociation rate coefficients to determine the modified Arrhenius form, and average energy change per dissociation to determine the novel two-parameter fit to $U$. This means that the goodness of fit demonstrated for the nonequilibrium test set is predictive. Predictive capability of this accuracy is a strong indicator that we are capturing the dominant physical mechanisms, instead of simply providing sufficient degrees of freedom to intersect the data. The Marrone-Treanor model is old and relatively simple, but it continues to be recommended for use by chemical kinetics authors [9, 43]. It has also been extended to exchange reactions by Knab et al. [9].
The modification we propose is consistent with qualitative and quantitative observations from the QCT data. Thus, the proposed modified Marrone-Treanor model is a simple, self-consistent description of dissociation at conditions characterized by $T$ and $T_{v}$.
(a) Dissociation rate coefficient, $T=T_{v}$
[…]
(c) Change in vibrational energy, $T=T_{v}$
[…]
Figure 7: Fitting to all QCT data for nitrogen dissociation data with new model: maximum relative error in rate coefficient is 22% and maximum error in vibrational energy change per dissociation is $0.24\,\mathrm{eV}$.
(a) Dissociation rate coefficient, $T=T_{v}$
[…]
(c) Change in vibrational energy, $T=T_{v}$
[…]
Figure 8: Fitting to all oxygen dissociation data with new model; effective dissociation with fixed $T_{D,s}$. The maximum relative error in rate is 15% and the maximum error in energy change is $0.18\,\mathrm{eV}$.
@@ 5 Non-Boltzmann Correction Factors
Section 2.2 introduced all analytical expressions needed for implementing the MMT model, which at its core consists of evaluating the dissociation rate coefficient with Eq. (10) and associated vibrational energy removed per dissociation according to Eq. (17). Thus the rate coefficient is evaluated as the product of the two-temperature expression $k_{d}^{\mathrm{arr}}(T)\,Z(T,T_{v})$ and a non-Boltzmann correction factor $f_{k}^{\mathrm{NB}}$. In similar fashion, the dissociating species’ energy removed per dissociation is evaluated as the product of Knab’s two-temperature formula for $\langle\varepsilon_{v,s}\rangle_{d}^{\mathrm{Knab}}(T,T_{v})$ and its own corresponding non-Boltzmann factor $f_{\varepsilon}^{\mathrm{NB}}$. Whereas the model’s dependence on $T$ and $T_{v}$ is justified by the analysis in Secs. 3 and 4, a more in-depth justification for including these correction factors was deferred to the present section.
@@
Section 1 referenced DMS [24, 14, 25, 26, 27] and state-to-state master equation studies [22] of coupled rovibrational excitation and dissociation in isothermal heat baths. Regardless of the mixture in question, all studies showed that after early transients the vibrational level populations of $\mathrm{N_{2}}$, $\mathrm{O_{2}}$ or $\mathrm{NO}$ adopted nearly time-invariant non-Boltzmann distributions as dissociation takes place. During this phase low-lying energy levels remain Boltzmann-populated at $T_{v}$, but the high-energy tail will be noticeably depleted [47, 48]. This depletion is a consequence of a dynamic balance established between preferential dissociation from high-$v$ levels and their comparatively slow replenishment through inelastic collisions. Such conditions have sometimes been referred to as the quasi-steady-state (QSS) dissociation regime [49] and the associated non-Boltzmann internal energy populations as QSS distributions. The analysis in Secs. 3 and 4 could not capture these non-Boltzmann depletion effects, since all vibrational populations were sampled from fixed Boltzmann distributions at a given $T_{v}$. Only dynamic calculation methods involving state-to-state master equations, or DMS can accurately establish the detailed dynamic balance between preferential dissociation and vibrational excitation that ultimately produces such QSS distributions [50]. Furthermore, vibrational population distributions with depleted high-energy tails are not exclusively a product of dissociation in isothermal heat baths. Similar non-Boltzmann distributions have also been observed in DMS calculations under adiabatic conditions [51] and downstream of normal shocks [52]. This implies that depletion effects are generally present in thermo-chemical nonequilibrium flow fields and must be taken into account as part of the MMT model formulation. From a macroscopic modeling viewpoint, when the diatomic species in a gas mixture adopt QSS vibrational distributions, this manifests in a reduction of dissociation rates (by factors $2-5$), as well as in a decrease of average vibrational energy removed per dissociation relative to values sampled from equivalent-temperature Boltzmann distributions.
@@
Figures. 9 and 10 illustrate the need to correct for such non-Boltzmann effects. Starting in Fig. 9(a) we plot the dissociation rate coefficient for reaction $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$. Black symbols represent the QCT-derived thermal equilibrium rate coefficient (where $T=T_{v}$), the continuous black line shows the corresponding Arrhenius fit (using Eq. (11) and parameters from Table 2). Red symbols represent the equivalent dissociation rate coefficient for quasi-steady-state (QSS) dissociation conditions, extracted from a series of isothermal heat bath DMS calculations [27] (where $T\gtrapprox T_{v}$). Within the temperature range shown, these QSS rate coefficients consistently remain below their corresponding thermal equilibrium counterparts by a near-constant factor. When the Arrhenius rate coefficient is re-scaled by $f_{k}^{\mathrm{NB}}=0.5$ (equivalent to Eq. (10) with $Z(T,T_{v}{=}T){=}1$) the dashed red line is obtained. The corresponding plot for $\mathrm{N_{2}}$ vibrational energy removed by dissociation is shown in Fig. 9(b). Black symbols again represent the QCT-derived values at thermal equilibrium, while the black continuous line shows the “Boltzmann” portion $\langle\varepsilon_{v,\mathrm{N_{2}}}\rangle_{d}^{\mathrm{Knab}}(T,T_{v}{=}T)$ evaluated from Knab’s formula with the MMT parameters $a_{U}$ and $U^{\star}$ from Table 2. Red symbols again represent the corresponding DMS-derived values extracted under QSS dissociation conditions. Analogous to what was observed for the rate coefficient, the $\mathrm{N_{2}}$ vibrational energy removed by dissociation at QSS conditions remains slightly below the corresponding thermal equilibrium QCT value over the entire temperature range shown. This behavior again suggests there exists a simple way of bringing the “Boltzmann” estimate of Eq. (16) closer in line with the DMS predictions. The dashed red line is obtained with Eq. (17), when $\langle\varepsilon_{v,\mathrm{N_{2}}}\rangle_{d}^{\mathrm{Knab}}(T,T_{v}{=}T)$ is multiplied by $f_{\varepsilon}^{\mathrm{NB}}=0.85$.
@@
Figures 9(c) and 9(d) present equivalent plots for reaction $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$, while analogous data for reactions $\mathrm{N_{2}+N\rightarrow 3N}$ and $\mathrm{O_{2}+O\rightarrow 3O}$ are compared in Fig. 10. All three reactions exhibit similar features to the first one: a dissociation rate coefficient about 50% smaller and roughly 15% less average vibrational energy removed per reaction in QSS than at thermal equilibrium. This consistent behavior makes it possible to capture the complex non-Boltzmann depletion effects in our two-temperature CFD model by employing the two simple correction factors. Of course, for every reaction the precise non-Boltzmann factors yielding the best possible fit between QCT and DMS data differ slightly and also vary somewhat with temperature. However, we choose to apply the same factors $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ to all four reactions so as to not introduce unnecessary complexity to the MMT model.
(a) Rate coefficient: $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$
(b) Vibrational energy removed per dissociation: $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$
(c) Rate coefficient: $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$
(d) Vibrational energy removed per dissociation: $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$
@@
(a) Rate coefficient: $\mathrm{N_{2}+N\rightarrow 3N}$
(b) Vibrational energy removed per dissociation: $\mathrm{N_{2}+N\rightarrow 3N}$
(c) Rate coefficient: $\mathrm{O_{2}+O\rightarrow 3O}$
(d) Vibrational energy removed per dissociation: $\mathrm{O_{2}+O\rightarrow 3O}$
@@ 6 Approach to Chemical Equilibrium
In Sec. 5 we motivated the use of non-Boltzmann correction factors as part of the MMT model by comparisons between QCT and DMS data. We found that values $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ produce very good $k_{d}$ and $\langle\varepsilon_{v,s}\rangle_{d}$ estimates for the dissociation reactions in air during the QSS dissociation phase. Thus, the MMT model is able to accurately reproduce the DMS-predicted dissociation- and vibrational energy removal rates experienced by the gas immediately downstream of a strong shock. However, employing constant non-Boltzmann factors becomes non-physical as the mixture approaches chemical equilibrium.
For illustration, consider the generic diatom dissociation-recombination reaction $\mathrm{AB}+\mathrm{M}\rightleftharpoons\mathrm{A}+\mathrm{B}+\mathrm{M}$. The forward and backward rates are evaluated according to Eq. (4) to be $\mathcal{R}_{r}^{f}=k_{r}^{f}\,\left(\rho_{\mathrm{AB}}/M_{\mathrm{AB}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ and $\mathcal{R}_{r}^{b}=k_{r}^{b}\,\left(\rho_{\mathrm{A}}/M_{\mathrm{A}}\right)\left(\rho_{\mathrm{B}}/M_{\mathrm{B}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ respectively. When the MMT model is employed, the forward (dissociation) rate coefficient $k_{r}^{f}$ is computed according to Eq (10). The backward (recombination) rate coefficient is evaluated using the same expression, but setting $Z(T,T)=1$ and invoking detailed balance to yield $k_{r}^{b}=f_{k}^{\mathrm{NB}}\,k_{d}^{\mathrm{arr}}(T)/K_{r}^{eq}(T)$. The net rate $\mathcal{R}_{r}=\mathcal{R}_{r}^{f}-\mathcal{R}_{r}^{b}$ then becomes:

### v2

@@ 4.4 Marrone-Treanor characteristic probability temperature
For each temperature condition in the equilibrium QCT dataset, we therefore determine the value of $U$ that best matches $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}(T,T_{\mathrm{v}}\!=\!T)$, via a Newton iteration on Eq. (16). We now have the dependence of $U$ on $T$, which we observe to be fit well by the linear relationship in inverse temperature space that is Eq. (15). We confirm this by plotting the aggregate results of the fitting procedure in Fig. 8 for the six reactions of Table 2. Nitrogen dissociation data in Fig. 8(a) were fit within the range $8\,000\,\mathrm{K}\leq T\leq 30\,000\,\mathrm{K}$, oxygen dissociation data in Fig. 8(b) were fit within the range $4\,000\,\mathrm{K}\leq T\leq 13\,000\,\mathrm{K}$.
By substituting Eq. (15) into Eq. (16) we may end up determining the model parameters $a_{U}$ and $U^{*}$ by simply fitting to $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}\,(T,T_{\mathrm{v}}=T)$ directly. In the end this means that the dependence of $U$ on $T$ is extracted entirely from the equilibrium vibrational energy change due to dissociation. The resulting values for $a_{U}$ and $U^{*}$ have been listed in Table 2. Note that these values for a given reaction differ between sub-tables 2(a) and 2(b) due to the precise values for $T_{\mathrm{D},s}$ employed.
@@
Figure 8: Dependence of Marrone-Treanor parameter $1/U$ on $1/T$, determined from vibrational energy loss per dissociation in equilibrium test set (symbols) and linear fit to Eq. (15) (lines).
@@ 4.5 Comparison of MMT model with QCT data
We now compare the original QCT data to the model prediction: in Fig. 9 for nitrogen dissociation and in Fig. 10 for oxygen dissociation. We find this modified Marrone-Treanor model to accurately fit all quantities of interest and for all dissociation reactions.
@@
It is worth recalling again that we have determined all model parameters using data in the equilibrium test set only: dissociation rate coefficients to determine the modified Arrhenius form, and average energy change per dissociation to determine the novel two-parameter fit to $U$. This means that the goodness of fit demonstrated for the nonequilibrium test set is predictive. Predictive capability of this accuracy is a strong indicator that we are capturing the dominant physical mechanisms, instead of simply providing sufficient degrees of freedom to intersect the data. The Marrone-Treanor model is old and relatively simple, but it continues to be recommended for use by chemical kinetics authors [9, 51]. It has also been extended to exchange reactions by Knab et al. [9].
The modification we propose is consistent with qualitative and quantitative observations from the QCT data. Thus, the proposed modified Marrone-Treanor model is a simple, self-consistent description of dissociation at conditions characterized by $T$ and $T_{\mathrm{v}}$.
(a) Dissociation rate coefficient, $T=T_{\mathrm{v}}$
(b) Dissociation rate coefficient, $T=20\,000\,\mathrm{K}$
(c) Change in vibrational energy, $T=T_{\mathrm{v}}$
(d) Change in vibrational energy, $T=20\,000\,\mathrm{K}$
Figure 9: Fitting to nitrogen dissociation QCT data. Maximum relative error in rate coefficient is 22% and maximum error in vibrational energy change is $0.24\,\mathrm{eV}$.
(a) Dissociation rate coefficient, $T=T_{\mathrm{v}}$
(b) Dissociation rate coefficient, $T=10\,000\,\mathrm{K}$
(c) Change in vibrational energy, $T=T_{\mathrm{v}}$
(d) Change in vibrational energy, $T=10\,000\,\mathrm{K}$
Figure 10: Fitting to oxygen dissociation QCT data. Maximum relative error in rate coefficient is 15% and maximum error in energy change is $0.18\,\mathrm{eV}$.
@@ 5 Non-Boltzmann correction factors
In Sec. 2.2 we introduced all analytical expressions needed for implementing the MMT model, which at its core consists of evaluating the dissociation rate coefficient with Eq. (10) and associated vibrational energy removed per dissociation according to Eq. (17). Thus the rate coefficient is evaluated as the product of the two-temperature expression $k^{\mathrm{Arr}}(T)\,Z(T,T_{\mathrm{v}})$ and a non-Boltzmann correction factor $f_{k}^{\mathrm{NB}}$. In a similar fashion, the dissociating species’ energy removed per dissociation is evaluated as the product of Knab’s two-temperature formula for $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}(T,T_{\mathrm{v}})$ and its own corresponding non-Boltzmann factor $f_{\varepsilon}^{\mathrm{NB}}$. Whereas the model’s dependence on $T$ and $T_{\mathrm{v}}$ is justified by the analysis in Secs. 3 and 4, a more in-depth justification for including these correction factors was deferred to the present section.
@@
In Sec. 1, we referenced DMS [30, 21, 31, 32, 34] and state-to-state master equation studies [26, 27, 28, 54] of coupled rovibrational excitation and dissociation of suddenly heated air species in isothermal heat baths. Regardless of the mixture in question, all these studies showed that after early transients the vibrational level populations of $\mathrm{N_{2}}$ and $\mathrm{O_{2}}$ adopted nearly time-invariant non-Boltzmann distributions as dissociation takes place with the respective $T_{\mathrm{v}}$ adopting a near-constant value slightly below the heat bath temperature $T$. During this phase low-lying energy levels remain Boltzmann-populated at $T_{\mathrm{v}}$, but the high-energy tail will be noticeably depleted [55, 56] below the expected Boltzmann populations. This depletion is a consequence of a dynamic balance established between preferential dissociation from high-$v$ levels and their comparatively slow replenishment through vibrational excitation, all the while recombination reaction rates which may also replenish the high-$v$ tail, are still negligible. Such conditions have sometimes been referred to as the QSS dissociation regime [53] and the associated non-Boltzmann internal energy populations as QSS distributions. A direct consequence of molecules adopting such depleted-tail non-Boltzmann distributions is that the macroscopic dissociation and vibrational energy removal rates during the QSS phase drop noticeably below their thermal equilibrium counterparts. The analysis in Secs. 3 and 4 could not capture these non-Boltzmann depletion effects, since all pre-collision vibrational populations were sampled from fixed Boltzmann distributions at a given $T_{\mathrm{v}}$. Only dynamic calculation methods involving state-to-state master equations, or DMS can accurately establish the detailed dynamic balance between preferential dissociation and vibrational excitation that ultimately produces such QSS distributions.
Methodologies for numerically estimating QSS dissociation rate coefficients can be found in literature dating back several decades. Even before the availability of the current ab initio-based QCT databases for air species, Gonzales and Varghese [57, 58] simulated nonequilibrium dissociation of various diatomic species in highly diluted isothermal heat baths using linearized master equations. They observed the gradual evolution from a vibrationally cold initial state toward the gas adopting its characteristic depleted-tail non-Boltzmann distribution and estimated the dissociation rate coefficient during the QSS phase. In subsequent work [59], they further showed that by including recombination reactions via detailed balance relations the gas eventually evolves past the QSS regime toward thermodynamic equilibrium characterized by the expected Boltzmann distribution at the heat bath temperature $T$. The more recent rovibrational-specific master equation and DMS studies [26, 27, 28, 54] for the $\mathrm{N_{2}}(v,j)-\mathrm{N}$ system all confirm the presence of these features.
Beyond this, vibrational population distributions with depleted high-energy tails are not exclusively a product of dissociation in isothermal heat baths. Similar non-Boltzmann distributions have also been observed in DMS calculations under adiabatic conditions [33] and downstream of normal shocks [60]. This implies that depletion effects are generally present in thermo-chemical nonequilibrium flow fields and must be taken into account as part of the MMT model formulation. From a macroscopic modeling viewpoint, when the diatomic species in a gas mixture adopt QSS vibrational distributions, this manifests in a reduction of dissociation rates (by factors 252-5), as well as in a decrease of roughly 10-20% in average vibrational energy removed per dissociation relative to values sampled from equivalent-temperature Boltzmann distributions.
@@
Figures 11 and 12 illustrate the need to correct for such non-Boltzmann effects. Starting in Fig. 11(a) we plot the dissociation rate coefficient for reaction $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$. Black symbols represent the QCT-derived thermal equilibrium rate coefficient (where $T=T_{\mathrm{v}}$), the continuous black line shows the corresponding Arrhenius fit (using Eq. (11) and parameters from Table 2). Red symbols represent the equivalent dissociation rate coefficient for quasi-steady-state (QSS) dissociation conditions, extracted from a series of isothermal heat bath DMS calculations [34] (where $T\gtrapprox T_{\mathrm{v}}$). Within the temperature range shown, these QSS rate coefficients consistently remain below their corresponding thermal equilibrium counterparts by a near-constant factor. When the Arrhenius rate coefficient is re-scaled by $f_{k}^{\mathrm{NB}}=0.5$ (equivalent to Eq. (10) with $Z(T,T_{\mathrm{v}}{=}T){=}1$) the dashed red line is obtained. The corresponding plot for $\mathrm{N_{2}}$ vibrational energy removed by dissociation is shown in Fig. 11(b). Black symbols again represent the QCT-derived values at thermal equilibrium, while the black continuous line shows the “Boltzmann” portion $\langle\varepsilon_{\mathrm{v},\mathrm{N_{2}}}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}(T,T_{\mathrm{v}}{=}T)$ evaluated from Knab’s formula with the MMT parameters $a_{U}$ and $U^{*}$ from Table 2. Red symbols again represent the corresponding DMS-derived values extracted under QSS dissociation conditions. Analogous to what was observed for the rate coefficient, the $\mathrm{N_{2}}$ vibrational energy removed by dissociation at QSS conditions remains slightly below the corresponding thermal equilibrium QCT value over the entire temperature range shown. This behavior again suggests there exists a simple way of bringing the “Boltzmann” estimate of Eq. (16) closer in line with the DMS predictions. The dashed red line is obtained with Eq. (17), when $\langle\varepsilon_{\mathrm{v},\mathrm{N_{2}}}\rangle_{\mathrm{diss}}^{\mathrm{Knab}}(T,T_{\mathrm{v}}{=}T)$ is multiplied by $f_{\varepsilon}^{\mathrm{NB}}=0.85$.
@@
Figures 11(c) and 11(d) present equivalent plots for reaction $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$, while analogous data for reactions $\mathrm{N_{2}+N\rightarrow 3N}$ and $\mathrm{O_{2}+O\rightarrow 3O}$ are compared in Fig. 12. All three reactions exhibit similar features to the first one: a dissociation rate coefficient about 50% smaller and roughly 15% less average vibrational energy removed per reaction in QSS than at thermal equilibrium. This consistent behavior makes it possible to capture the complex non-Boltzmann depletion effects in our two-temperature CFD model by employing the two simple correction factors. Of course, for every reaction the precise non-Boltzmann factors yielding the best possible fit between QCT and DMS data differ slightly and also vary somewhat with temperature. However, we choose to apply the same factors $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ to all four reactions so as to not introduce unnecessary complexity to the MMT model.
(a) Rate coefficient for $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$
(b) Vibrational energy removed per dissociation for $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$
(c) Rate coefficient for $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$
(d) Vibrational energy removed per dissociation for $\mathrm{2O_{2}\rightarrow 2O+O_{2}}$
@@
(a) Rate coefficient for $\mathrm{N_{2}+N\rightarrow 3N}$
(b) Vibrational energy removed per dissociation for $\mathrm{N_{2}+N\rightarrow 3N}$
(c) Rate coefficient for $\mathrm{O_{2}+O\rightarrow 3O}$
(d) Vibrational energy removed per dissociation for $\mathrm{O_{2}+O\rightarrow 3O}$
@@ 6 Approach to chemical equilibrium
In Sec. 5 we motivated the use of non-Boltzmann correction factors as part of the MMT model by comparisons between QCT and DMS data. We found that values $f_{k}^{\mathrm{NB}}=0.5$ and $f_{\varepsilon}^{\mathrm{NB}}=0.85$ produce very good $k_{\mathrm{diss}}$ and $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$ estimates for the major nitrogen and oxygen dissociation reactions during the QSS dissociation phase. Thus, the MMT model is able to accurately reproduce the DMS-predicted rates of dissociation and vibrational energy removal experienced by the gas immediately downstream of a strong shock. However, blindly employing that same pair of constant non-Boltzmann factors becomes non-physical as the dissociating mixture approaches chemical equilibrium.
For illustration, consider the generic diatom dissociation-recombination reaction $\mathrm{AB}+\mathrm{M}\rightleftharpoons\mathrm{A}+\mathrm{B}+\mathrm{M}$. The forward and backward rates are evaluated according to Eq. (4) to be $\mathcal{R}_{\mathrm{diss}}=k_{\mathrm{diss}}\,\left(\rho_{\mathrm{AB}}/M_{\mathrm{AB}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ and $\mathcal{R}_{\mathrm{rec}}=k_{\mathrm{rec}}\,\left(\rho_{\mathrm{A}}/M_{\mathrm{A}}\right)\left(\rho_{\mathrm{B}}/M_{\mathrm{B}}\right)\left(\rho_{\mathrm{M}}/M_{\mathrm{M}}\right)$ respectively. In a two-temperature model the forward (dissociation) rate coefficient at an arbitrary nonequilibrium condition $T\neq T_{\mathrm{v}}$ becomes $k_{\mathrm{diss}}=k_{\mathrm{diss}}(T,T_{\mathrm{v}})$ and in MMT specifically it is computed from Eq (10). But calculating the backward (recombination) rate coefficient $k_{\mathrm{rec}}$ is not as straightforward, especially in a nonequilibrium setting. Despite recent QCT studies [61] on the available ab initio PESs, at present we are unaware of any general and reliable method to calculate the recombination rates directly from first principles. The main impediments, as we see them, are the added complexity to sampling initial parameters for the recombining three-body collisions (as opposed to dissociating two-body collisions), as well as the lack of a fully developed kinetic theory to calculate a three-body collision rate.
The most commonly used approach is to instead tie the recombination rate to its corresponding dissociation rate using detailed balance relations. Such relations would be expressed at the kinetic scale by linking together cross sections for forward and backward reactions, but a reasonable simplification at the fluid scale is to assume Maxwellian velocity distributions at translational temperature $T$ for all collision partners. In this limit it becomes possible to formulate detailed balance relations in terms of rate coefficients and the equilibrium constant of a given reaction. This approach has for instance been used in master equation studies [26, 27, 28] to enforce detailed balance relations in terms of state-specific dissociation-recombination rate coefficient pairs and drive the simulated mixture to thermodynamic equilibrium. Beyond this, in two-temperature models such as Park [53] or MMT, it is implicitly assumed that the low-lying vibrational levels of diatomic species are populated according to Boltzmann distributions at temperature $T_{\mathrm{v}}$. Therefore, the rate coefficients of forward and backward reactions at thermal equilibrium must naturally be evaluated at the single equilibrium temperature $T_{\mathrm{v}}=T$.
In the present discussion we proceed as follows. First, recall the detailed balance relation that holds at thermodynamic equilibrium:
$K_{\mathrm{c}}^{\mathrm{eq}}(T)=\left[\frac{\left(\rho_{\mathrm{A}}/M_{\mathrm{A}}\right)\left(\rho_{\mathrm{B}}/M_{\mathrm{B}}\right)}{\left(\rho_{\mathrm{AB}}/M_{\mathrm{AB}}\right)}\right]^{\star}=\frac{k_{\mathrm{diss}}^{\star}(T)}{k_{\mathrm{rec}}^{\star}(T)}.$ (26)
Here $K_{\mathrm{c}}^{\mathrm{eq}}(T)$ is the equilibrium constant for the generic dissociation-recombination reaction and the $\star$” superscripts indicate that all quantities take on their respective thermodynamic equilibrium values. As implied by the first equality in Eq. (26), the equilibrium constant at temperature $T$ is equivalent to the equilibrium product-to-reactant molar concentration ratio, which is the one that minimizes the Gibbs free energy. In practice, the $K_{\mathrm{c}}^{\mathrm{eq}}$ are typically calculated from thermodynamic databases, such as the NASA Lewis polynomial fits [62] and are themselves independent of the reaction rates. But at thermodynamic equilibrium the second equality in Eq. (26) must also hold. It stems from of the fact that at this specific condition the net reaction rate must be zero: